OpenStudy (anonymous):

The function f is given by f (x) = e^(x–11) –8. (a) Find f –1(x). (b) Write down the domain of f –1(x).

OpenStudy (anonymous):

-1 = inverse...

OpenStudy (anonymous):

Kinda confused

OpenStudy (anonymous):

sorry! not sure.. this might help tho, http://www.zweigmedia.com/RealWorld/calctopic1/inverses.html

OpenStudy (zzr0ck3r):

\[y=e^{x-11}-8\\\text{we want to solve for x in terms of y}\\\text{add 8 to both sides}\\y+8=e^{x-11}\\\text{take natural log of both sides}\\ln(y+8)=ln(e^{x-11})=(x-11)ln*(e)=x-11\\so\\ln(y+8)=x-11\\\text{add 11 to both sides}\\ln(y+8)+11=x\\\text{now switch the x's with y's so that we have a function of x}\\y=ln(x+8)-11\]

OpenStudy (zzr0ck3r):

@trace60 let me know if this makes sense.

OpenStudy (agent0smith):

n/m, @zzr0ck3r answered it in the exact same way I would :)

OpenStudy (anonymous):

But wait whats the domain? =S

OpenStudy (zzr0ck3r):

the domain depends on the definition of the function, but I'm guessing it is all real numbers for codomain. having said that, is there any value of x we cant put in our final equation?

OpenStudy (agent0smith):

Domain of original function = range of inverse function Range of original function = domain of inverse function So start with the original function - find it's domain and range.

OpenStudy (anonymous):

No, there are no values for x and thank you

OpenStudy (agent0smith):

What's the range of f (x) = e^(x–11) –8 ? this will be your domain of the inverse.

OpenStudy (zzr0ck3r):

The range of f(x) is f(x)>=-8, so if it were going to have an inverse it must be onto and 1-1 its codomain. so the domain of the inverse, while having no problems on the real line, will have a restricted domain.

OpenStudy (anonymous):

Would a table of values be necessary in the finding of the domain?

OpenStudy (agent0smith):

No, what's the minimum possible value of f (x) = e^(x–11) –8? think of it like this... if e^(x-11) is really really really small, what's f(x) equal to?

OpenStudy (anonymous):

Finding the domain and range has never been my cup of tea so like =/ ...still not with you on that one

OpenStudy (agent0smith):

If e^(x-11) is zero, then what's f (x) = e^(x–11) –8 equal to?

OpenStudy (agent0smith):

Note that e^(x-11) can never be a negative number, it can only be almost zero, or positive.

OpenStudy (agent0smith):

If e^(x-11) is zero, then f(x) = 0 - 8

OpenStudy (anonymous):

So then f(x) =0? *squints eyes*

OpenStudy (agent0smith):

Try again :P

OpenStudy (anonymous):

8

OpenStudy (agent0smith):

Try one more time...

OpenStudy (anonymous):

-.-

OpenStudy (agent0smith):

0-8 is not positive 8

OpenStudy (agent0smith):

So -8 is the minimum possible value for the range. What's the max possible value for f (x) = e^(x–11) –8 ?

OpenStudy (agent0smith):

Or, is there one?

OpenStudy (agent0smith):

So there's no limit, so the range is from y=-8 to infinity. That becomes the domain of your inverse, so the domain of x is...?

OpenStudy (agent0smith):

Oops i deleted that. So your range for y is from -8 to infinity This becomes your domain for the inverse function.

OpenStudy (anonymous):

Thank you